OpenStudy (anonymous):

If A^2 = [ 7 10 ] find A. Help Please. [ 15 22 ]

OpenStudy (anonymous):

[ 7 10 ] [ 15 22 ] <---- matrix A^2

OpenStudy (anonymous):

help please?...

OpenStudy (ikram002p):

Trace =7+22=29 det=7*22-10*15=139 let S^2=det and t^2=Trace+2s S=+-sqrt (139)=+13,-13 t=+-sqrt (29+13)=+sqrt 42,-sqrt42 t=+-sqrt(29-13)=4,-4 so for A^2=[ a b] [c d] A=1/t [a+s b] [c d+s] so u wud have more than one matrix

OpenStudy (amoodarya):

\[A=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\\\] find A^2 parametric then solve the system of equation

OpenStudy (ikram002p):

i think ull have 16 matrix of A :o

OpenStudy (anonymous):

@ikram002p about the det=7*22-10*15=139... its 154-150 right?... it is equal to 139?..

OpenStudy (ikram002p):

ohh sry i made a typo , yess 154-150=4 :o so i need to check every thing nw :(

OpenStudy (anonymous):

it's alright...

OpenStudy (ikram002p):

Trace =7+22=29 det=7*22-10*15=4 let S^2=det and t^2=Trace+2s S=+-sqrt (4)=+2,-2 t=+-sqrt (29+2*2)=+sqrt 33,-sqrt33 t=+-sqrt(29-2*2)=+5,-5

OpenStudy (ikram002p):

let me see one solution so for A^2=[ a b] [c d] A=1/t [a+s b] [c d+s] when t=5 s=2 A=1/5 [7+2 10] [15 22+2] A=[9/5 2] [3 24/5] lets check from A^2

OpenStudy (anonymous):

umm.. isn't t=5 and s=-2 would be better? because it will become, 1/5 [ 7-2 10 ] [ 15 22-2 ] ... 1/5 [ 5 10 ] [ 15 20 ] therefore, [ 1 2 ] A = [ 3 4 ] umm, check me if I'm wrong...

OpenStudy (ikram002p):

ur right but as i said there is 16 matrix of A that gives u A^2 as u posted :D so u got one solution of 16

OpenStudy (anonymous):

oh. ^^ ok thanks. ummm, I know I'm asking too much but does this procedure also applicable if I would be looking for the root of a 3x3 matrix?

OpenStudy (ikram002p):

NO , this procedure only work for 2×2 matrix

OpenStudy (anonymous):

so... looking for the root of a 3x3 matrix would be more complex?

OpenStudy (ikram002p):

here u can check quickly from the root http://ncalculators.com/matrix/2x2-matrix-multiplication-calculator.htm

OpenStudy (ikram002p):

no its not complecated give a matrix A = [x1 x2 x3] [x4 x5 x6] [x7 x8 x9] then find A^2 so u wud have a linear system of equation hmmm u cud solve it for 9 variable or use a program that solve linear equations its easy lol

OpenStudy (anonymous):

umm, can you give an example if it's ok with you?

OpenStudy (ikram002p):

u give me an example ^^ and ill solve it for u

OpenStudy (anonymous):

ok... hmm. wait.

OpenStudy (anonymous):

[ 2 1 2 ] A = [ 3 -3 -4 ] [ 4 -1 2 ]

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

oh that must be A^2..

OpenStudy (ikram002p):

ok u creat this example or u find it in a book ? cuz if u creat it it might not work or well have complex number got me ?

OpenStudy (anonymous):

oh... I just create it... ^^ ' sorry

OpenStudy (ikram002p):

huh lol humm wud u like to work on it ?

OpenStudy (anonymous):

[ 1 2 1 ] A^2 = [ 2 1 1 ] [ 2 1 2 ] I think smaller elements would be simple (I hope) how about this?

OpenStudy (anonymous):

I just created it..

OpenStudy (ikram002p):

do u have matlab program ? cuz if u use matlab there is a code that gives u the the root in few sec

OpenStudy (ikram002p):

ok... lets work on it lol

OpenStudy (ikram002p):

A = [x1 x2 x3] [x4 x5 x6] [x7 x8 x9] find A^2

OpenStudy (anonymous):

umm, no. don't know matlab program. sorry..

OpenStudy (anonymous):

ok ^^

OpenStudy (ikram002p):

A^2= [x1 x2 x3] [x1 x2 x3] [x4 x5 x6] [x4 x5 x6] [x7 x8 x9] [x7 x8 x9]

OpenStudy (ikram002p):

can u do the multiple thing ??

OpenStudy (anonymous):

yes

OpenStudy (ikram002p):

ok wat u r waiting for lol do it :P

OpenStudy (anonymous):

[ 30 36 42 ] A^2 = [ 66 81 96 ] [ 102 126 150 ] am I correct? umm, by the way, what's the "x" infront of the numbers?

OpenStudy (ikram002p):

lol no ! not the matrix u gave the matrix i gave u

OpenStudy (anonymous):

yes yes, that's the 123456789 matrix..

OpenStudy (anonymous):

(_ _) sorry if I'm too bad at this...

OpenStudy (ikram002p):

A^2= [x1 x2 x3] [x1 x2 x3] [x4 x5 x6] [x4 x5 x6] [x7 x8 x9] [x7 x8 x9] =[x1*x1+x2*x4+x3*x7 x1*x2+x2*x5+x3*x8 x1*x3+x2*x6+x3*x9] [x4*x1+x5*x4+x6*x7 x4*x2+x5*x5+x6*x8 x4*x3+x5*x6+x6*x9] [x7*x1+x8*x4+x9*x7 x7*x2+x8*x5+x9*x8 x7*x3+x8*x6+x9*x9 ] ok ?

OpenStudy (anonymous):

yes, that's what I did.

OpenStudy (ikram002p):

lol ok lets continue :P

OpenStudy (anonymous):

what's the "x" infront of the numbers?

OpenStudy (ikram002p):

[x1*x1+x2*x4+x3*x7 x1*x2+x2*x5+x3*x8 x1*x3+x2*x6+x3*x9] [1 2 1 ] [x4*x1+x5*x4+x6*x7 x4*x2+x5*x5+x6*x8 x4*x3+x5*x6+x6*x9]= [2 1 1] [x7*x1+x8*x4+x9*x7 x7*x2+x8*x5+x9*x8 x7*x3+x8*x6+x9*x9 ] [2 1 2] x1,......,x9 are variables u can put a,b,c..... instead ok ??

OpenStudy (anonymous):

ok

OpenStudy (ikram002p):

x1*x1+x2*x4+x3*x7 =1 x1*x2+x2*x5+x3*x8=2 x1*x3+x2*x6+x3*x9=1 x4*x1+x5*x4+x6*x7=2 x4*x2+x5*x5+x6*x8=1 x4*x3+x5*x6+x6*x9=1 x7*x1+x8*x4+x9*x7=2 x7*x2+x8*x5+x9*x8=1 x7*x3+x8*x6+x9*x9=2 so u have 9 equation solve to find \(x_1,x_2,x_3,x_4,x_5,x_6,x_7,x_8,x_9\)

OpenStudy (ikram002p):

nw solving those equation lol u need a program or if u really good solve them :D

OpenStudy (anonymous):

O.O .....

OpenStudy (ikram002p):

huh :P in wat class u r ?

OpenStudy (anonymous):

what.. class? mmm, advanced algebra?

OpenStudy (ikram002p):

ok i guess u dnt need to solve qs like this anf u had to, theyll give u simple numbers or somthin u just need to knw how to make equations not solving them ok ?

OpenStudy (anonymous):

ok. ^^

OpenStudy (anonymous):

is there an equation for the 2x2 matrix above?

OpenStudy (anonymous):

i guess this would be my last question, I'd already took too much of your time...

OpenStudy (ikram002p):

yess there is just use @amoodarya method

OpenStudy (ikram002p):

\(A=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\\\) \(A^2=\left[\begin{matrix}a×a+b×c & a×b+b×d \\ c×a+d×c & c×b+d×d\end{matrix}\right]\\\) \(A^2=\left[\begin{matrix}7 & 10 \\ 15 & 22\end{matrix}\right]\\\) \(\left[\begin{matrix}7 & 10 \\ 15 & 22\end{matrix}\right]=\left[\begin{matrix}a×a+b×c & a×b+b×d \\ c×a+d×c & c×b+d×d\end{matrix}\right]\) \(a^2+bc=7\) \(ab+bd=10\) \(ac+dc=15\) \(cb+d^2=22\)

OpenStudy (ikram002p):

ok good luck g2g ^^

OpenStudy (anonymous):

ok, thanks. ^^

OpenStudy (ikram002p):

np